# NuPlanQA: A Large-Scale Dataset and Benchmark for Multi-View Driving Scene Understanding in Multi-Modal Large Language Models

### v1

@@ 3 Experiments
VideoLLaMA27B [10] 50.2 55.8 92.2 66.1 66.9 77.0 63.6 69.2 68.6 81.7 80.9 77.1 70.8
Qwen2.5-VL7B [2] 49.8 68.7 64.1 60.9 45.3 12.2 24.5 27.3 22.5 35.1 8.3 22.0 36.7
LLaVA-OV7B [19] 49.8 88.9 95.8 78.2 76.2 76.5 71.4 74.7 72.8 73.3 81.9 76.0 76.3
LLaVA-NV7B [54] 40.9 44.2 68.8 51.3 53.0 46.9 38.2 46.0 55.0 62.9 67.3 61.7 53.0
InternVL-1.520B [9] 36.5 67.7 71.9 58.7 43.3 46.9 36.8 42.3 53.4 51.0 53.8 52.7 51.2
LLaVA-NV32B [54] 44.8 71.0 91.1 69.0 65.7 70.9 60.9 65.8 70.2 65.8 74.9 70.3 68.4
@@
VideoLLaMA27B [10] 47.8 61.8 92.7 67.4 65.2 77.0 60.0 67.4 67.5 82.7 80.4 76.9 70.6
Qwen2.5-VL7B [2] 50.7 67.3 58.3 58.8 43.6 10.2 18.6 24.1 22.0 36.6 7.0 21.9 34.9
LLaVA-OV7B [19] 49.8 92.6 96.4 65.9 72.4 74.5 67.3 71.4 69.1 70.8 79.9 73.3 70.2
LLaVA-NV7B [54] 46.8 35.0 57.3 46.4 49.2 46.4 35.9 43.8 52.4 51.0 66.3 56.6 48.9
InternVL-1.520B [9] 35.6 64.5 67.2 55.8 51.9 44.4 35.9 44.1 60.2 52.0 54.3 55.5 51.8
LLaVA-NV32B [54] 43.6 70.5 89.1 67.7 65.2 70.4 51.8 62.5 70.2 66.3 81.9 72.8 67.7
@@ 3.1 Settings
Comparison of Recent MLLMs. We evaluate video LLMs, including both commercial and open-source models, on our NuPlanQA-Eval benchmark to assess their inherent performance in traffic scene analysis. Our evaluation is conducted in a zero-shot setting. Due to the absence of publicly available multi-view MLLMs incorporating multi-view images, BEV-LLM serves as the multi-view trained baseline. For other MLLMs, we use structured multi-view images as inputs by arranging them into a single image, marking the corresponding camera view for reference. Each model’s default settings are used to adjust the frame size accordingly. Since our benchmark consists of multiple-choice questions with four answer options, we adopt accuracy as the evaluation metric, where random guessing yields a baseline accuracy of 25%.

### v2

@@ 3 Experiments
VideoLLaMA27B [10] 54.2 59.4 93.8 69.1 70.7 78.6 66.4 71.9 72.3 82.7 81.4 78.8 73.3
Qwen2.5-VL7B [2] 51.7 71.9 69.3 64.3 49.2 23.0 30.9 34.4 30.4 39.6 10.1 26.7 41.8
LLaVA-OV7B [19] 53.2 89.4 96.4 79.7 77.9 77.6 73.2 74.3 75.4 76.2 82.9 77.5 77.8
LLaVA-NV7B [54] 44.3 46.5 72.4 54.4 57.5 52.0 44.1 51.2 57.6 65.8 71.4 64.9 56.8
InternVL-1.520B [9] 40.9 70.5 75.5 62.3 48.6 50.0 44.5 47.7 57.1 55.0 60.8 57.6 55.9
LLaVA-NV32B [54] 47.8 74.7 93.8 72.1 69.1 74.0 63.6 68.9 72.8 69.3 76.9 73.0 71.3
@@
VideoLLaMA27B [10] 50.2 64.1 93.8 69.4 71.3 79.6 64.1 71.7 69.6 84.2 81.4 78.4 73.2
Qwen2.5-VL7B [2] 51.2 70.5 62.5 61.4 47.0 16.8 23.2 29.0 27.2 40.6 12.1 26.6 39.0
LLaVA-OV7B [19] 52.2 93.5 96.4 80.7 72.9 74.0 68.6 71.8 72.3 72.3 79.9 74.8 75.8
LLaVA-NV7B [54] 48.8 41.5 63.5 51.3 51.9 50.5 41.8 48.1 56.5 60.4 69.8 62.2 53.9
InternVL-1.520B [9] 39.4 68.7 72.4 60.2 54.7 48.5 41.4 48.2 62.3 58.9 58.8 60.0 56.1
LLaVA-NV32B [54] 46.8 73.3 91.7 70.6 70.7 71.9 58.2 66.9 71.2 68.8 82.9 74.3 70.6
@@ 3.1 Settings
Comparison of Recent MLLMs. We evaluate video LLMs, including both commercial and open-source models, on our NuPlanQA-Eval benchmark to assess their inherent performance in traffic scene analysis. Due to the absence of publicly available multi-view MLLMs incorporating multi-view images, BEV-LLM serves as the multi-view trained baseline. For other MLLMs, we use structured multi-view images as inputs by arranging them into a single image, marking the corresponding camera view for reference. Each model’s default settings are used to adjust the frame size accordingly. Since our benchmark consists of multiple-choice questions with four answer options, we adopt accuracy as the evaluation metric, where random guessing yields a baseline accuracy of 25%.